# Motion Manifold Flow Primitives for Language-Guided Trajectory Generation

### v1

@@ 4 Language-based Robot Motion Generation
Robust text-conditional latent flow learning: The latent space flow-based model consists of two neural networks. One is a text embedding neural network $h_{\theta}$ that maps $c\mapsto\tau=h_{\theta}(c)\in{\cal T}$, where ${\cal T}=\mathbb{R}^{p}$ is called a text embedding space. The other is a neural network vector field $v(\cdot;\theta)$ in the latent space that maps $(t,z,\tau)\mapsto v_{t}(z,\tau;\theta)\in T_{z}{\cal Z}$, where $T_{z}{\cal Z}$ is the tangent space of ${\cal Z}$ at $z$. We train these two networks simultaneously with the following regularized flow matching loss:
$\min_{\theta}\frac{1}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}\Big{(}\mathbb{E}_{t,z}\big{[}\|v_{t}(z,h_{\theta}(c_{ij});\theta)-u_{t}(z|z_{i})\|^{2}\big{]}+\frac{\gamma}{K}\sum_{k=1}^{K}\|h_{\theta}(c_{ij})-h_{\theta}(\tilde{c}^{k}_{ij})\|^{2}\Big{)}.$ (4)
@@
In the expectation of the first term, $t,z$ are sampled from ${\cal U}[0,1]$, $p_{t}(z|z_{i})$, respectively, and $p_{t}(z|z_{i})$ is defined as the OT Gaussian path and $u_{t}(z|z_{i})$ is derived from it (see section 3.2). The second term, multiplied by a positive weight $\gamma$, is added to ensure robustness against diverse text variations. Here, $\tilde{c}^{k}_{ij}$ are Sentence-BERT encoding vectors that have meanings similar to those of $c_{ij}$, which are generated via the Large Language Model ChatGPT.
@@
Using the trained text embedding model $h_{\theta}$, vector field $v_{\theta}$, and decoder $f_{\theta}$, we can generate, given a free-form text input, diverse trajectories by the following procedure (see Figure 2): (i) encode the text input using the Sentence-BERT to $c$, (ii) embed $c$ to a text embedding vector $\tau=h_{\theta}(c)$, (iii) solve an ODE $\dot{z}=v_{t}(z,\tau;\theta)$ from $t=0$ to $t=1$ with initial samples $z_{0}\sim p_{0}(z)={\cal N}(z|0,I)$ and obtain $z_{1}\in{\cal Z}$, and (iv) map the ODE solutions $z_{1}$ to trajectories $x=f_{\theta}(z_{1})$. We refer our entire framework to as Motion Manifold Flow Primitives (MMFP).
@@ 5 Experiments
In this section, we evaluate our method, the MMFP, mainly compared to (i) Denoising Diffusion Probabilistic Models (DDPM) [9] trained in the trajectory space ${\cal X}$ – when ${\cal X}$ is ${\rm SE}(3)^{T}$, we train DDPM in local coordinates –, (ii) Flow Matching (FM) [12] trained in the trajectory space ${\cal X}$ – when ${\cal X}$ is ${\rm SE}(3)^{T}$, we use the Riemannian Flow Matching (RFM) [13] – (iii) Task-Conditional Variational Autoencoder (TCVAE) with Gaussian prior [16], and (iv) Motion Manifold Primitives with Gaussian mixture prior (MMP) [15]. MMFP trained without the regularization term (i.e., the second term in (4)) in the latent flow learning will be denoted by MMFP w/o reg. We compare models for 6-DoF SE(3) pouring trajectory generation and 7-DoF robot arm waving motion generation tasks. Training details are available in the Supplementary Material.
@@
Evaluation metrics: To measure the similarity between the generated trajectories given a text input and demonstration trajectories annotated with that text, we use the Maximum Mean Discrepancy (MMD) [59]; the lower, the better. Texts are annotated to each trajectory in three different levels (e.g. see Figure 3 Left) and the MMD metrics are measured separately for each text input, and then averaged across text descriptions at the same level. To evaluate the robustness to text variations, we also report robust MMD metrics. These are computed with unseen unbiased text inputs generated by ChatGPT. Lastly, to evaluate whether the model generates accurate motions that perform the task described in the given text, we train a trajectory classifier, and use it to report the motion accuracy; the higher, the better. More details are available in the Supplementary Material.
@@ 5.1 Text-based pouring motion generation
DDPM [9] 0.398 0.484 1.306 0.400 0.486 1.304 50.5 20.0 10.2
RFM [13] 0.411 0.431 0.778 0.413 0.425 1.127 87.0 36.4 30.8
TCVAE [16] 0.117 0.211 0.824 0.131 0.191 1.041 52.0 25.4 15.3
MMP [15] 0.045 0.115 0.950 0.055 0.112 0.970 47.5 19.6 9.3
@@
Figure 5: The procedure of a feasible joint space trajectory generation.
@@
Now, when provided with SE(3) trajectories of a bottle sampled using our MMFP, we propose a method to quickly find feasible joint space trajectories for an arbitrary 2D position of the cup; see Figure 5 for the overview. To begin, for a given SE(3) trajectory of a bottle, we leverage symmetries in the pouring task to augment it into diverse SE(3) trajectories. This augmentation involves three types of transformations, as illustrated in Figure 5 (Trajectory augmentation via symmetry transformations): (1) rotation of the bottle around the cup axis by $\theta_{b}\in[-30\degree,30\degree]$, (2) rotation of the bottle frame by $\theta_{g}\in[-60\degree,60\degree]$ that changes the grasping pose – where the robot reaches and grasps the bottle in the green axis direction –, and (3) flipping the bottle with respect to the cup (i.e., rotating it by $180\degree$). Each generated trajectory undergoes a total of $50=5\times 5\times 2$ transformations. This involves discretizing $[-30\degree,30\degree]$ and $[-60\degree,60\degree]$ into 5 points with equal-size intervals for $\theta_{b}$ and $\theta_{g}$, respectively. Additionally, multiples of 2 are considered for flipping. We generate five trajectories for a single text input, resulting in a total of 250 candidate SE(3) trajectories that require verification for the existence of an IK solution. We train and use a neural IK classifier to boost the speed of finding feasible joint trajectories. An example SE(3) trajectory whose IK solution exists is shown in Figure 5 (Feasible joint space trajectory); other examples can be found in the Supplementary Material.
@@
Table 2: Real robot success rates given language guidance.
We conduct real robot experiments with two scenarios. First, we put an obstacle blocking the middle path, and ask the robot “Give me some water, and pour it from the very left/right side.”, so that it pours water while turning away to avoid collision; see Figure 6 (Upper). Second, we ask the robot “Give me some wine/water.”, and see if it shows clockwise rotation behavior of the bottle at the end of the pouring; see Figure 6 (Lower). We generate three joint space trajectories for each text input, and measure the success rates. Success is judged by human observation. Table 2 shows the success rates [^2] Trajectories from DDPM and RFM are too jerky to be deployed in the real robot.. MMFP outperforms TCVAE and MMP in all experiments. We note that, in EXP 3, the success rate of MMFP is 2/3. Judging by human observation, one trajectory is somewhat ambiguous to determine whether the robot is pouring water or pouring wine. Figure 6 shows examples where MMFP successfully avoids the obstacle and shows the correct pouring style.
@@ 5.2 Text-based 7-DoF waving motion generation
In this section, we train text-based 7-DoF waving motion generation models, where the dataset is obtained from human demonstrations. The demonstrator is instructed to hold and move the robot arm so that the robot arm mimics waving motions in five different viewing directions (i.e., very left, left, front, right, and very right) and in three different styles (i.e., very big, big, small). We extract 7-DoF joint space trajectories, and the trajectory lengths are pre-processed to ensure $T=720$. We encode two motions for each setting, resulting in a total of $N=30$ trajectories. For each trajectory $x_{i}\in{\cal X}$, we provide three text annotations, denoted by $\{c_{ij}\}_{j=1}^{M}$ for $M=3$, as shown in Figure 7.
@@
DDPM [9] 0.425 0.542 0.831 0.427 0.542 0.831 18.4 29.3 4.1
FM [12] 0.211 0.368 0.646 0.215 0.384 0.671 99.6 88.7 91.1
TCVAE [16] 0.269 0.581 0.833 0.294 0.606 0.865 25.6 57.3 10.7
MMP [15] 0.013 0.369 0.772 0.016 0.355 0.750 19.2 30.7 6.5
@@
6 Limitation
Learned motion manifolds effectively interpolate given demonstration trajectories, but struggle to extrapolate and generate more varied motions. For instance, in our pouring task, the demonstrations are restricted to turning left or right, so the model is not able to produce motions turning upwards or downwards. Collecting more trajectory data could solve this problem, but data collection is costly. Ideally, a solution that bypasses the need for additional data would be preferable.

### v2

@@ 4.2 Robust Text-conditional Latent Flow Learning
The latent space flow-based model consists of two neural networks. One is a text embedding neural network $h_{\theta}$ that maps $c\mapsto\tau=h_{\theta}(c)\in{\cal T}$, where ${\cal T}=\mathbb{R}^{p}$ is called a text embedding space. The other is a neural network vector field $v(\cdot;\theta)$ in the latent space that maps $(s,z,\tau)\mapsto v_{s}(z,\tau;\theta)\in T_{z}{\cal Z}$, where $T_{z}{\cal Z}$ is the tangent space of ${\cal Z}$ at $z$. We train these two networks simultaneously with the following regularized flow matching loss:
$\min_{\theta}\frac{1}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}\Big{(}$ $\mathbb{E}_{s,z}\big{[}\|v_{s}(z,h_{\theta}(c_{ij});\theta)-u_{s}(z|z_{i})\|^{2}\big{]}$
$+\frac{\gamma}{K}\sum_{k=1}^{K}\|h_{\theta}(c_{ij})-h_{\theta}(\tilde{c}^{k}_{ij})\|^{2}\Big{)}.$ (8)
@@
In the expectation of the first term, $s,z$ are sampled from ${\cal U}[0,1]$, $p_{s}(z|z_{i})$, respectively, and $p_{s}(z|z_{i})$ is defined as the OT Gaussian path and $u_{s}(z|z_{i})$ is derived from it (see the preliminary section on flow matching). The second term, multiplied by a positive weight $\gamma$, is added to ensure robustness against diverse text variations. Here, $\tilde{c}^{k}_{ij}$ are Sentence-BERT encoding vectors that have meanings similar to those of $c_{ij}$, which are generated via the Large Language Model ChatGPT.
@@
Using the trained text embedding model $h_{\theta}$, vector field $v_{\theta}$, and decoder $f_{\theta}$, we can generate, given a free-form text input, diverse trajectories by the following procedure (see Figure 2): (i) encode the text input using the Sentence-BERT to $c$, (ii) embed $c$ to a text embedding vector $\tau=h_{\theta}(c)$, (iii) solve an ODE $z^{\prime}=v_{s}(z,\tau;\theta)$ from $s=0$ to $s=1$ with initial samples $z_{0}\sim p_{0}(z)={\cal N}(z|0,I)$ and obtain $z_{1}\in{\cal Z}$ – where we use Euler method with $ds=0.05$ –, and (iv) map the ODE solutions $z_{1}$ to trajectories $x=f_{\theta}(z_{1})$. We call this framework Motion Manifold Flow Primitives (MMFP).
We note that our MMFP, which solves the ODE in the latent space to obtain a single latent value and maps it to the entire trajectory $z\mapsto x=(q_{1},\ldots,q_{T})$, fundamentally differs from conventional latent ODE methods that model and map the latent trajectory $(z_{1},\ldots,z_{T})$ to the trajectory $x=(q_{1},\ldots,q_{T})$ (Rubanova et al., 2019; Zhi et al., 2022).
Lastly, one may ask a reasonable question: Why not use diffusion models (Song et al., 2020b) in the latent space instead of flow models, which are known for their ability to fit complex conditional distributions? We note that flow-based models trained via flow matching – rather than the conventional maximum likelihood method – can be understood as deterministic versions of diffusion models (Lipman et al., 2022). In our trajectory generation tasks, empirical results show that latent flow models outperform diffusion models (see the experiment section on the comparison of latent diffusion and flow models). This performance gap is likely due to the relatively small size of our trajectory dataset.
@@ 5 Experiments
In this section, we evaluate our method, the MMFP, mainly compared to (i) Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) trained in the trajectory space ${\cal X}$ – when ${\cal X}$ is ${\rm SE}(3)^{T}$, we train DDPM in local coordinates –, (ii) Flow Matching (FM) (Lipman et al., 2022) trained in the trajectory space ${\cal X}$ – when ${\cal X}$ is ${\rm SE}(3)^{T}$, we use the Riemannian Flow Matching (RFM) (Chen and Lipman, 2023) – (iii) Task-Conditional Variational Autoencoder (TCVAE) with Gaussian prior (Noseworthy et al., 2020), and (iv) Motion Manifold Primitives with Gaussian mixture prior (MMP) (Lee et al., 2023). MMFP trained without the regularization term (i.e., the second term in (4.2)) in the latent flow learning will be denoted by MMFP w/o reg.
For MMFP, we choose the latent space dimension $m=3$ for ${\cal Z}=\mathbb{R}^{m}$ and text embedding dimension $p=3$ for ${\cal T}=\mathbb{R}^{p}$. For TCVAE and MMP, with thorough tuning, we set the latent space dimension to be $3$ or $4$ and text embedding dimension to be $64$. We compare models for 6-DoF SE(3) pouring trajectory generation and 7-DoF robot arm waving motion generation tasks. Throughout, we use fully-connected neural networks with ELU activation functions.
Evaluation metrics: The primary objective is to generate correct motions corresponding to the given text inputs. To evaluate whether the model generates accurate motions that perform the task described in the given text, we train a trajectory classifier and use it to report the motion accuracy; the higher, the better. Additionally, the model should be able to generate not just a single trajectory but diverse trajectories, imitated from the given demonstration dataset, depending on the text input. For example, see Figure 6 and 12. We use the Maximum Mean Discrepancy (MMD) (Gretton et al., 2012) – which measures the distance between two probability distributions, computed from their samples – to measure the similarity between the set of generated trajectories given a text input and the set of demonstration trajectories annotated with that text; the lower, the better. A formal mathematical expression of the MMD metric is provided below: let $\{x_{i}\}_{i=1}^{N}$ and $\{y_{j}\}_{j=1}^{M}$ be the samples of $p(x)$ and $p(y)$ in the metric space ${\cal X}$ with a distance metric denoted by $d_{\cal X}(\cdot,\cdot)$. First, we define a symmetric positive kernel $k(x,y):=\exp(-\frac{d^{2}_{\cal X}(x,y)}{h})$, where $h:=\sum_{i,j=1}^{N+M}\frac{d^{2}_{\cal X}(z_{i},z_{j})}{(N+M)(N+M-1)}$ for $z_{i}\in\{x_{1},\ldots,x_{N},y_{1},\ldots,y_{M}\}$. Then the MMD metric ${\rm MMD}(p(x),p(y))$ is computed as
$\scriptsize\frac{1}{N^{2}}\sum_{i,j=1}^{N}k(x_{i},x_{j})+\frac{1}{M^{2}}\sum_{i,j=1}^{M}k(y_{i},y_{j})-\frac{2}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}k(x_{i},y_{j}).$ (9)
@@
Texts are annotated to each trajectory at multiple different levels (e.g., see Figure 5 Left and Figure 10), and the MMD metrics are measured separately for each level text input, then averaged across text descriptions at the same level. Evaluation metrics are computed using both seen and unseen texts in training. First, accuracy and MMD metrics are measured using the seen texts. Then, to evaluate the robustness to text variations, we also report robust MMD metrics, which are computed with unseen unbiased text inputs generated by ChatGPT.
We begin this section by comparing latent diffusion and flow models using two-dimensional collision-free trajectory generation tasks. Additionally, we present experiments with higher-dimensional configuration spaces, the text-based SE(3) pouring motion generation and 7-DoF waving motion generation tasks.
5.1 Comparison of Latent Diffusion and Flow Models with 2D Trajectory Generation
In this section, we provide deeper insights into latent flow matching algorithms by comparing flows trained with the Optimal Transport (OT) path and various diffusion-based paths. Recall that we first train a motion manifold of trajectories, followed by training a flow model in the latent coordinate space. The flow matching algorithm includes a key design choice: selecting the conditional probability paths $p_{s}(x|x_{1})$ (see the preliminary section on the flow matching). This section investigates how different path choices affect the outcomes. For illustration purposes, we use a simple example: a text-based 2D trajectory generation task.
Lipman et al. (2022) demonstrated that flow matching in certain cases can recover diffusion processes. Diffusion models (Song et al., 2020a; Ho et al., 2020) gradually corrupt data by adding noise, which leads the data to approach pure noise over time. These processes can be represented as Gaussian conditional probability paths, $p_{s}(x|x_{1})={\cal N}(x|\mu_{s}(x_{1}),\sigma^{2}_{s}(x_{1})I)$, where the forward diffusion process corresponds to $s=1\to s=0$. Specific choices of mean $\mu_{s}(x_{1})$ and variance $\sigma_{s}(x_{1})$ determine the nature of the path.
For example, Variance Exploding (VE) path has the form
$p_{s}(x|x_{1})={\cal N}(x\>|x_{1},\sigma^{2}_{1-s}I),$ (10)
where $\sigma_{s}$ is an increasing function, $\sigma_{0}=0$ and $\sigma_{1}\gg 1$. The Variance Preserving (VP) diffusion paths has the form
$p_{s}(x|x_{1})={\cal N}(x\>|\alpha_{1-s}x_{1},(1-\alpha^{2}_{1-s})I),$ (11)
[…]
Lipman et al. (2022) also showed that these diffusion paths align with the vector fields used in deterministic probability flow for score-based models (Song et al., 2020b). Compared to diffusion paths, they found that the OT Gaussian path results in smoother ODE trajectories. This makes it easier for neural network vector fields to fit the data, improving accuracy and reducing the number of function evaluations required during sampling.
Figure 3: Demonstration trajectories with multiple text annotations. Each trajectory is assigned with two text labels, the level 1 and level 2 texts.
Table 1: MMFPs trained with OT and various diffusion paths. The same motion manifold (i.e., an autoencoder) is employed.
MMD ($\downarrow$) Accuracy ($\uparrow$)
level 1 level 2
path
task
Diffusion-VE
0.075 0.007 100 94.6
Diffusion-VP-1
0.073 0.003 100 95.0
Diffusion-VP-2
0.030 0.055 94.3 81.0
Optimal Transport
0.025 0.004 100 99.8
We reproduce and reinterpret these findings – originally obtained with image datasets – using our own 2D trajectory dataset with text annotations (shown in Figure 3) and extend them with a novel interpretation in the context of MMFP. In our dataset, there are 20 demonstration trajectories that do not collide with the wall, all starting from the same point and ending at a common goal point (the origin). There are four distinct passages to the origin, with five trajectories passing through each passage (upper, lower, right, and left). Each trajectory has a length of $T=201$, and the trajectory space is ${\cal X}=\mathbb{R}^{2T}$. For each trajectory $x_{i}\in{\cal X}$, we assign two text annotations: level 1 (the same for all trajectories, “Go to the origin”) and level 2 (providing more specific instructions, shared among groups of five trajectories).
We use the same motion manifold (an autoencoder) and train latent vector fields with different Gaussian probability paths: variance exploding path with $\sigma_{1-s}=3(1-s)$ denoted by Diffusion-VE, variance preserving paths with $\alpha_{1-s}=e^{-5(1-s)^{2}}$ denoted by Diffusion-VP-1 and $\alpha_{1-s}=e^{-0.5(1-s)^{2}}$ denoted by Diffusion-VP-2, and the optimal transport path.
Table 1 shows Diffusion-VE, Diffusion-VP-1, Diffusion-VP-2, and Optimal Transport, the last of which is utilized in our MMFP throughout the paper. We note that Optimal Transport path consistently yields the best results. Figure 4 illustrates the evolution of generated trajectories following each of the latent vector fields, described by $z^{\prime}_{s}=v_{s}(z_{s},\tau)$, where at each $z_{s}$ is mapped to a 2D trajectory through the decoder function. The trajectories displayed on the far right represent the final output samples. Optimal Transport is shown to provide the best outcomes.
Figure 4: Evolution of generated trajectories (from left to right) follows each of the latent vector fields trained with different Gaussian paths, including both diffusion paths and the Optimal Transport path, with the trajectories on the far right representing the final output samples.
Unlike image-based generative models trained with large datasets, the limited data in our case amplifies the impact of the chosen probability path. While in image models the primary advantage of using the Optimal Transport path is faster sampling, here it improves generative performance by learning a model with a smoother velocity field, which proves to be crucial for generalization with small trajectory datasets. Therefore, the latent flow models of our MMFP presented in the subsequent section are always trained with Optimal Transport paths.
@@ 5.2 Text-based pouring motion generation
DDPM (Ho et al., 2020) 0.398 0.484 1.306 0.400 0.486 1.304 50.5 20.0 10.2
RFM (Chen and Lipman, 2023) 0.411 0.431 0.778 0.413 0.425 1.127 87.0 36.4 30.8
TCVAE (Noseworthy et al., 2020) 0.117 0.211 0.824 0.131 0.191 1.041 52.0 25.4 15.3
MMP (Lee et al., 2023) 0.045 0.115 0.950 0.055 0.112 0.970 47.5 19.6 9.3
@@
To understand why TCVAE and MMP fail, it’s essential to understand the conditional AutoEncoder (cAE) architecture. In a cAE, the decoder takes a latent value and a text embedding vector together as inputs to generate a trajectory as the output. However, the single decoder network in a cAE struggles to fit a complex, non-smooth function like the one in this example, where the text-conditional trajectory distribution changes dramatically across level 1, 2, and 3 texts. On the other hand, our latent flow-based models can capture such complex distributions because they smoothly transform the distributions not all at once but gradually via the ODE.
We provide a visual analysis of the latent distributions of the SE(3) pouring dataset in our MMFP. Trajectory data themselves are very high-dimensional and hence difficult to visualize. Fortunately, since our motion manifold models use three-dimensional latent spaces, the distributions of the data in the latent space can be visualized. Additionally, the text embedding space also has a dimension of three, so text embedding vectors can be visualized as well. Figure 7 shows the latent distributions of both the trajectory data and the text embeddings.
Figure 8 highlights the multi-modality of the text-conditioned latent distribution and, more importantly, the dramatic change in the distribution as the text inputs vary. In the figure, moving from left to right corresponds to scenarios where level 1 to level 3 texts are given as inputs. When a level 1 text is provided, the distribution must fit two modalities, and the volume of the distribution is the largest. In contrast, when level 2 or level 3 texts are provided, there is only one modality. Note that the support of the distribution varies significantly from level 1 to level 3, becoming much smaller at level 3. This implies that a latent text-conditioned distribution should capture such dramatic changes, fitting a complex, non-smooth function. Conditional autoencoder models struggle to fit such a complex function, while our latent flow models are capable of doing so.
Figure 7: Upper: SE(3) pouring trajectory data encoded in the three-dimensional latent space by our motion encoder $g_{\theta}$. Lower: Texts used in training encoded in the three-dimensional text embedding space by our text embedding model $h_{\theta}$.
Figure 8: Sampled latent points by our trained latent flow model in MMFP. Left: Sampled latent points given the level 1 text. Middle: Sampled latent point given a level 2 text. Right: Sampled latent points given a level 3 text.
@@
Table 3: Real robot success rates given language guidance.
@@
We conduct real robot experiments with two scenarios. First, we put an obstacle blocking the middle path, and ask the robot “Give me some water, and pour it from the very left/right side.”, so that it pours water while turning away to avoid collision, without relying on vision systems but by correctly understanding human instructions; see Figure 9 (Upper). Second, we ask the robot “Give me some wine/water.”, and see if it shows clockwise rotation behavior of the bottle at the end of the pouring; see Figure 9 (Lower). Given generated SE(3) trajectories of the bottle, we compute joint space trajectories by solving the inverse kinematics. Since this process can be computationally intensive, we train and use a neural IK classifier for fast and feasible joint trajectory computation; details can be found in Appendix A. We generate three joint space trajectories for each text input, and measure the success rates. Success is judged by human observation. Table 3 shows the success rates (trajectories from DDPM and RFM are not implemented because they are too jerky to be deployed in the real robot). MMFP outperforms TCVAE and MMP in all experiments. We note that, in EXP 3, the success rate of MMFP is 2/3. Judging by human observation, one trajectory is somewhat ambiguous to determine whether the robot is pouring water or pouring wine. Figure 9 shows examples where MMFP successfully understands human instructions, thereby avoiding the obstacle and demonstrating the correct pouring style.
Regarding the obstacle avoidance experiments, one may ask, what if it is difficult for the user to instruct the avoidance direction? We note that if a vision system capable of recognizing obstacles and performing collision checking is available, then our framework can indeed be equipped with an autonomous obstacle avoidance mechanism. This can be achieved by searching for a proper latent value that leads to a collision-free trajectory, i.e., finding one that is collision-free within the motion manifold. Specifically, given multiple latent samples, we can use a decoder function $f_{\theta}$ to map them to trajectories, and then check whether the trajectories are collision-free. If the obstacle blocks the trajectory pouring from the center, trajectories that avoid it by pouring from the left or right will be sampled. However, if there are very strict constraints, such as multiple obstacles blocking all paths, a solution may not exist. Intuitively, learning a larger motion manifold that contains a greater diversity of motions increases the likelihood of finding a solution. Therefore, ensuring diversity in the demonstration data, which limits the diversity of the motion manifold, although costly, is crucial for improving the adaptability of the system.
@@ 5.3 Text-based 7-DoF waving motion generation
In this section, we train text-based 7-DoF waving motion generation models, where the dataset is obtained from human demonstrations. The demonstrator is instructed to hold and move the robot arm so that the robot arm mimics waving motions in five different viewing directions (i.e., very left, left, front, right, and very right) and in three different styles (i.e., very big, big, small). We extract 7-DoF joint space trajectories, and the trajectory lengths are pre-processed to ensure $T=720$. Thus, ${\cal Q}=\mathbb{R}^{7}$, and the trajectory space is denoted as ${\cal X}=\mathbb{R}^{7\times T}$. We encode two motions for each setting, resulting in a total of $N=30$ trajectories. For each trajectory $x_{i}\in{\cal X}$, we provide three text annotations, denoted by $\{c_{ij}\}_{j=1}^{M}$ for $M=3$, as shown in Figure 10. Thus, multiple trajectories can be associated with a single text label; for example, all 30 trajectories are labeled with the level 1 text, ‘Wave your hand.’. In the following, we compare our MMFP with the FM, TCVAE and MMP, both quantitatively and qualitatively.
@@
[…]
@@
DDPM (Ho et al., 2020) 0.425 0.542 0.831 0.427 0.542 0.831 18.4 29.3 4.1
FM (Lipman et al., 2022) 0.211 0.368 0.646 0.215 0.384 0.671 99.6 88.7 91.1
TCVAE (Noseworthy et al., 2020) 0.269 0.581 0.833 0.294 0.606 0.865 25.6 57.3 10.7
MMP (Lee et al., 2023) 0.013 0.369 0.772 0.016 0.355 0.750 19.2 30.7 6.5
@@
Figure 13: Real robot waving motion generation by MMFP.